FUTURE CONTRACTS

Dudley Savings Bank wishes to take a position in Treasury bond futures contracts, which currently have a quote of 111 − 100. Dudley Savings thinks interest rates will go down over the period of investment. The face value of the bond underlying the futures contract is $100,000.
 
a. Should the bank go long or short on the futures contracts?
b. Given your answer to part (a), calculate the net profit to Dudley Savings Bank if the price of the futures contracts increases to 111 − 240. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))
c. Given your answer to part (a), calculate the net profit to Dudley Savings Bank if the price of the futures contracts decreases to 110 − 300. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))
Answer #1

a). If interest rates decrease over the period of investment, Treasury bond prices will increase. Thus, Dudley Savings Bank should take a long position in the futures contracts on the Treasury bonds. As T-bond prices go up, so will T-bond futures prices.

b). Given a long position:

Purchase price of futures = 111 − 100

= [111 + (10/32)]% × $100,000 = 111.3125% x $100,000 = $111,312.50

Sale price of futures = 111 − 240

= [111 + (24/32)]% × $100,000 = 111.75% x $100,000 = $111,750

Net profit = Sale price of futures − Purchase price of futures

= $111,750 − $111,312.50 = $437.50

c). Given a long position:

Purchase price of futures = 111 − 100

= [111 + (10/32)]% × $100,000 = 111.3125% x $100,000 = $111,312.50

Sale price of futures = 110 − 300

= [110 + (30/32)]% × $100,000 = 110.9375% x $100,000 = $110,937.50

Net profit = Sale price of futures − Purchase price of futures

= $110,937.50 − $111,312.50 = -$375


answered by: Blacketi